An electric motor develops 5.0-kW at 2000 rpm.The motor delivers its power through a spur gear 20-cm in diameter.If two teeth of the gear system to torque to another gear at the same instant, find the force exerted by each gear tooth.


1
Expert's answer
2020-06-01T14:27:17-0400

Power "(P)" , Speed rpm "(N)" , Torque "(T)"


"p = \frac {2\pi NT}{60}"


"T = \frac{5000 \times 60}{2\pi\times2000}"


"T = 23.87 Nm"


the force exerted by each gear tooth "(F)"


Diameter "(D) = 20 cm = 0.2 m"


"T = F \times \frac{D}{2}"


"F = \frac{23.87 \times 2}{0.2} = 238.7N"